# 0-supple.tex

### v1

@@ Abstract
As tensors become widespread in modern data analysis, Tucker low-rank Principal Component Analysis (PCA) has become essential for dimensionality reduction and structural discovery in tensor datasets. Motivated by the common scenario where large-scale tensors are distributed across diverse geographic locations, this paper investigates tensor PCA within a distributed framework where direct data pooling is impractical.
@@ 1 Introduction
In recent years, the prevalence of large-scale tensor datasets across diverse fields has garnered significant attention. Tucker low-rank tensor Principal Component Analysis (PCA) and Singular Value Decomposition (SVD) have become indispensable for unsupervised learning and dimension reduction, proving crucial in deriving valuable insights across various applications. These techniques are extensively utilized in recommendation systems (Karatzoglou et al., 2010), natural image and video processing (Gatto et al., 2021), neuroimaging analysis (Li et al., 2018), and health care systems (Ren et al., 2023), among others.
@@
In the second scenario with heterogeneity in Section 3, tensors observed on different machines are allowed to be generated from different underlying models that share some common principal components. Our focus is to improve estimation across all machines handling heterogeneous data. We formally define the partition of common and individual components for each tensor and develop a generalized distributed Tensor PCA method to efficiently estimate these shared components and also identify and extract unique components specific to each tensor. This approach allows us to accommodate the distinct characteristics of each dataset within the collective framework.
The third scenario with heterogeneity in Section 4 focuses on another setting of heterogeneous distributed learning, aiming at enhancing the estimation accuracy of a tensor at a specific location by intelligently integrating abundant data from other locations. We develop a novel transfer learning algorithm based on a weighted averaging scheme that involves calibrating the influence of data based on their noise levels and relevance, thereby improving the precision of the target site’s tensor estimation.
We have established comprehensive statistical error bounds for our proposed estimators across different scenarios. In Section 2.3, our analysis involves a detailed decomposition and quantification of bias and variance terms. We demonstrate that our estimator achieves the optimal minimax rate when the signal-to-noise ratio (SNR) exceeds a certain threshold, thus equating the performance of the pooled estimator that aggregates tensors directly (further details in Section 2.1). Essentially, our distributed algorithm attains the highest possible estimation accuracy achievable in a non-distributed setting.
In the heterogeneous scenario detailed in Section 3.2, we show that our estimator for common components matches the performance rate of the homogeneous scenario, while the estimators for individual components maintain rates consistent with individual tensor PCA. This underscores the effectiveness of our method in simultaneously learning shared and unique tensor structures. Importantly, even when pooling of tensors is possible, creating an efficient estimator through direct aggregation is challenging due to data heterogeneity—a fact corroborated by our numerical studies where our distributed methods surpass the pooled estimator in various heterogeneous configurations.
Furthermore, we provide theoretical guarantees for our estimator designed to utilize knowledge from multiple distributed sites, optimizing the weight allocations in our algorithm to enhance estimation accuracy beyond what is achievable with the target dataset alone. In addition, we derive the asymptotic distribution of the distance between our proposed estimator and the true values in Section B of the supplementary material, facilitating statistical inference and enabling the construction of confidence regions for the singular subspaces.
@@
Our work, while related to distributed matrix PCA, tackles the more complex issue of distributed tensor PCA, which is inherently more challenging both practically and technically. Unlike matrices, tensors often involve three or more modes, increasing their dimensionality and complicating their analysis. Whereas matrices might be addressed through convex methods like nuclear norm penalization, tensors require nonconvex formulations due to their more complex structure, involving multiple low-rank components and a core tensor. This introduces significant challenges in algorithm design and the establishment of theoretical guarantees, including statistical convergence and inference, which are critical for distributed tensor PCA. The contributions of our work are summarized as follows.
@@
  • Theoretical Contributions: We establish statistical guarantees for our distributed methods, demonstrating that they achieve a sharp statistical error rate that aligns with the minimax optimal rate possible in a non-distributed setting. Additionally, we calculate the asymptotic distribution of the distance between our estimator and the true values, as detailed in Section B of the supplementary material. This calculation aids in statistical inference and supports the construction of confidence regions for the singular subspaces. These contributions represent a theoretical advancement in utilizing aggregated common components, a finding that has not been previously documented in tensor analysis literature and enhances the existing computational approaches to distributed Tucker decomposition.
@@ 1.1 Related works
Motivated by modern scientific research, tensor decompositions have been actively studied in machine learning, electrical engineering, and statistics. Despite the well-established theory for low-rank decomposition of matrices, tensors present unique challenges. There are several notions of low-rankness in tensors (Kolda and Bader, 2009), including the most popular CANDECOMP/PARAFAC (CP) low-rankness and multilinear/Tucker low-rankness. Richard and Montanari (2014), Hopkins et al. (2015), and Perry et al. (2020) considered a rank-1 spiked statistical tensor PCA model, a special case of CP decomposition, and proposed various methods, including tensor unfolding and sum of square optimization. Zhang and Xia (2018) introduced a general framework of tensor SVD based on Tucker decomposition and established statistical and computational limits, followed by Xia et al. (2022) who further studied inference for tensor PCA. Zhang and Han (2019) considered sparse tensor SVD where the loading matrices are assumed to be sparse. Wang and Li (2020) studied the decomposition of a higher-order tensor with binary entries. Chen et al. (2024) further introduced a general framework of semi-parametric tensor factor models based on tensor PCA with auxiliary covariate information.
@@
To handle the challenges posed by massive and decentralized data, there has been a significant amount of recent literature developing distributed estimation or inference techniques for a variety of statistical problems (Li et al., 2013; Chen and Xie, 2014; Garber et al., 2017; Lee et al., 2017; Shi et al., 2018; Jordan et al., 2019; Fan et al., 2019; Volgushev et al., 2019; Chen et al., 2019; Li et al., 2020; Yu et al., 2020; Zhao et al., 2020; Tan et al., 2022; Luo et al., 2022; Lv and Lian, 2022; Yu et al., 2022; Chen et al., 2022). We refer readers to Gao et al. (2022) for a comprehensive review of distributed learning.
Among these works, our paper is most closely related to the literature on distributed matrix PCA. Fan et al. (2019) notably proposed a convenient one-shot approach that computes the top-$K$-dim eigenspace of the covariance matrix on each local machine and aggregates them on a central machine. They established a rigorous statistical guarantee showing that this approach achieves a sharp error rate as long as the sample size on each local machine is sufficiently large. An alternative multi-round method was simultaneously developed by Garber et al. (2017), which mainly estimates the first eigenspace by leveraging shift-and-invert preconditioning. Subsequently, Chen et al. (2022) proposed an improved multi-round algorithm for estimating the top-$K$-dim eigenspace that enjoys a fast convergence rate under weaker restrictions compared to Fan et al. (2019). Other works focused on providing more general statistical error bounds (Zheng and Tang, 2022) or improving the communication efficiency (Huang et al., 2021; Charisopoulos et al., 2021).
@@
The structure of this paper is as follows: Section 2 introduces our distributed tensor PCA algorithm for homogeneous settings and establishes its statistical error rates. Section 3 adapts the algorithm for heterogeneous environments and includes theoretical analysis. Section 4 explores the extension to transfer learning. Section 5 presents simulations and real data analyses to evaluate the performance of our methods. Some commonly used notations in this paper are defined in Section A of the supplementary material. Section B, notably, presents the asymptotic distribution of our proposed estimator, featuring distributed inference. All technical proofs and additional numerical experiments are presented in Sections C and D of the supplementary material, respectively.
@@ 2 Estimation for Homogeneous Tensors
In this section, we present the distributed tensor PCA algorithm for the homogeneous setting where each machine observes a tensor generated from the same underlying model. We first formulate the problem setup and illustrate the pooled estimator as a benchmark in Section 2.1, followed by the intuition behind its construction detailed in Section 2.2. Finally, Section 2.3 establishes theoretical guarantees on the statistical error rate for our algorithm.
@@ 2.1 Problem Setup and the Pooled Estimator
If the tensors $\{\mathcal{T}_{\ell}\}$ were allowed to be pooled onto a central machine, a standard way for estimating the singular space $\mathrm{Col}(U_{j})$ would have been to conduct a Tucker decomposition on the averaged tensor $\overline{\mathcal{T}}=\frac{1}{L}\sum\limits_{\ell=1}^{L}\mathcal{T}_{\ell}=\mathcal{T}^{*}+\overline{\mathcal{Z}}$, where $\overline{\mathcal{Z}}=\frac{1}{L}\sum\limits_{\ell=1}^{L}\mathcal{Z}_{\ell}$ is a tensor with i.i.d. $\mathcal{N}(0,\frac{\sigma^{2}}{L})$ entries. Denote the estimator obtained through this method as $\widehat{U}_{\mathrm{pooled},j}$, $j\in[J]$, which serves as a benchmark against which we evaluate the effectiveness of our approach in the distributed environment.
@@ 2.3 Theoretical Guarantee
Assume that there exist constants $C_{1},c_{1},C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma\lambda_{\min}^{-1}$ for any $j\in[J]$ and $\ell\in[L]$. If $p_{j}\asymp p$ for all $j$, $L\lesssim p^{c_{3}}$ for some $c_{3}>0$, $p\gtrsim r^{J-1}$, $\kappa_{0}=O(1)$, and $\lambda_{\min}/\sigma\gtrsim\sqrt{pr}$, then we have
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}=O_{\mathbb{P}}\bigg{(}\frac{\sigma}{\lambda_{\min}}\sqrt{\frac{pr}{L}}+\frac{pr\sigma^{2}}{\lambda_{\min}^{2}}\bigg{)},$ (3)
where $\{\widehat{U}_{j}\}$ is the output of Algorithm 1.
@@
where $G_{1}=\mathcal{M}_{1}(\mathcal{G})$, $Z_{1,\ell}=\mathcal{M}_{1}(\mathcal{Z}_{\ell})$, $U_{1\perp}$ is the orthogonal complement of $U_{1}$, and $R_{1,\ell}$ is a remainder term. When the signal-to-noise ratio (SNR) satisfies $\lambda_{\min}/\sigma\gtrsim\sqrt{pr}$, the Frobenius norm of the first two mean-zero terms on the RHS of (4) is of the order $O\left(\sqrt{pr}\sigma\lambda_{\min}^{-1}\right)$, and the remainder term $R_{1,\ell}$ has a higher order $O\big{(}pr\sigma^{2}\lambda_{\min}^{-2}\big{)}$. By averaging the projection matrices on all machines, the order of the first two terms can be reduced to $O\big{(}\sqrt{pr}\sigma\lambda_{\min}^{-1}L^{-1/2}\big{)}$, while that of $R_{1,\ell}$ does not change, leading to the error rate in (3). When the SNR is sufficiently large with respect to $L$, the first term in (3) dominates the second one, which leads to the following corollary.
@@
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}=O_{\mathbb{P}}\bigg{(}\frac{\sigma}{\lambda_{\min}}\sqrt{\frac{pr}{L}}\bigg{)}.$ (5)
@@ 3 Estimation for Heterogeneous Tensors
$\mathcal{T}_{\ell}^{*}=\mathcal{G}_{\ell}\times_{1}[U_{1}\;V_{1,\ell}]\times_{2}[U_{2}\;V_{2,\ell}]\cdots\times_{J}[U_{J}\;V_{J,\ell}],\quad U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}},V_{j,\ell}\in\mathbb{O}^{p_{j}\times r_{j,V,\ell}},\mathcal{G}_{\ell}\in\mathbb{R}^{r_{1,\ell}\times r_{2,\ell}\cdots\times r_{J,\ell}},$ (7)
where $r_{j,\ell}=r_{j,U}+r_{j,V,\ell}$ and $U_{j}^{\top}V_{j,\ell}=0$. For each $j$, the component $U_{j}$ spans a common singular subspace shared by all tensors $\{\mathcal{T}_{\ell}\}$, while $V_{j,\ell}$ spans an individual subspace specific to each tensor. Moreover, the core $\mathcal{G}_{\ell}$ is allowed to be different for different $\ell$. The goal is to estimate the shared singular subspace and the individual subspace separately for $j\in[J]$ and $\ell\in[L]$. Note that in this section, we assume the noise level $\sigma^{2}$ is identical on all machines for clear presentation. The heterogeneity on $\sigma^{2}$ will be further investigated in Section 4.
@@
where $\Lambda_{j,\ell}$ is a diagonal matrix with decreasing singular values for all $j\in[J]$, $\ell\in[L]$. The condition (8) is consistent with the Identification Condition (Assumption 1) in Chen et al. (2024), which ensures the identifiability of $\big{(}\mathcal{G}_{\ell}$, $[U_{1}\;V_{1,\ell}],\dots,[U_{J}\;V_{J,\ell}]\big{)}$ in model (7) when the singular values are distinct.
@@
To illustrate why (8) ensures the identifiability, consider the mode-$j$ matricization of $\widetilde{\mathcal{G}}$ under model (1), i.e., $\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}\big{)}=O_{j}^{\top}G_{j}\big{(}\bigotimes\limits_{k\neq j}O_{k}\big{)}$, where $G_{j}=\mathcal{M}_{j}(\mathcal{G})$. There exists a unique $O_{j}\in\mathbb{O}^{r_{j}\times r_{j}}$, the left singular value matrix of $G_{j}$, such that $\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}\big{)}\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}\big{)}^{\top}=O_{j}^{\top}G_{j}G_{j}^{\top}O_{j}$ is a diagonal matrix with distinct decreasing singular values, and thus $\widetilde{U}_{j}=U_{j}O_{j}$ is uniquely determined.

### v3

# Distributed Tensor Principal Component Analysis with Data Heterogeneity

@@ Abstract
As tensors become widespread in modern data analysis, Tucker low-rank Principal Component Analysis (PCA) has become essential for dimensionality reduction and structural discovery in tensor datasets. Motivated by the common scenario where large-scale tensors are distributed across diverse geographic locations, this paper investigates tensor PCA within a distributed framework where direct data pooling is theoretically suboptimal or practically infeasible.
@@ 1 Introduction
In recent years, the prevalence of large-scale tensor datasets across diverse fields has garnered significant attention. Tucker low-rank tensor Principal Component Analysis (PCA) and Singular Value Decomposition (SVD) have become indispensable for unsupervised learning and dimension reduction, proving crucial in deriving valuable insights across various applications. These techniques are extensively utilized in recommendation systems (Karatzoglou et al., 2010), biomedicine (Wen et al., 2024), natural image and video processing (Gatto et al., 2021), and health care systems (Ren et al., 2023), among others.
@@
In the second scenario in Section 3, tensors observed on different machines are allowed to be generated from different underlying models that share common leading principal components. Our focus is to improve estimation across all machines handling heterogeneous data. We formally define the partition of the common and individual components for each tensor and develop a generalized distributed Tensor PCA method to efficiently estimate these shared components and also identify and extract unique components specific to each tensor. This approach allows us to accommodate the distinct characteristics of each dataset within the collective framework.
The third scenario in Section 4 focuses on another setting of heterogeneous distributed learning, aiming at enhancing the estimation accuracy of a tensor at a specific location by intelligently integrating abundant data from other locations. We develop a novel knowledge transfer algorithm based on a weighted averaging scheme that involves calibrating the influence of data based on their noise levels and relevance, thereby improving the precision of the target site’s tensor estimation.
We have established comprehensive statistical error results for our proposed estimators across different scenarios. In Section 2.3, our analysis involves a detailed decomposition and quantification of the bias and variance terms. We demonstrate that our estimator achieves the optimal minimax rate when the signal-to-noise ratio (SNR) exceeds a certain threshold, thus equating the performance of the pooled estimator that aggregates tensors directly (further details in Section 2.1). Essentially, our distributed algorithm attains the highest possible estimation accuracy achievable in a non-distributed setting.
In the heterogeneous scenario detailed in Section 3.2, we show that our estimator for common components matches the performance of the homogeneous scenario, while the estimators for individual components maintain rates consistent with individual tensor PCA. This underscores the effectiveness of our method in simultaneously learning shared and unique tensor structures. Importantly, even when pooling of tensors is possible, creating an efficient estimator through direct aggregation is challenging due to data heterogeneity—a fact corroborated by our numerical studies where our distributed methods surpass the pooled estimator in various heterogeneous configurations.
Furthermore, we provide theoretical guarantees for our knowledge-transfer estimator, optimizing the weight allocations in our algorithm to enhance estimation accuracy beyond what is achievable with the target dataset alone. In addition, we derive the asymptotic distribution of the distance between our proposed estimator and the truth in Section E of the supplementary material, facilitating statistical inference and enabling the construction of confidence regions for the singular subspaces.
@@
Our work, while related to distributed matrix PCA (e.g., Fan et al., 2019; Chen et al., 2022), tackles the more complex issue of distributed tensor PCA, which is inherently more challenging both practically and technically. Unlike matrices, tensors often involve three or more modes, increasing their dimensionality and complicating their analysis. The estimators of tensor PCA are often calculated from optimization algorithms iteratively over multiple modes of low-rank components and a core tensor, whereas the estimator of matrix PCA is based on non-iterative schemes. Due to the intricate statistical dependencies involved in iterative optimization algorithms, the establishment of theoretical guarantees for the distributed estimators in tensor PCA, including statistical convergence and inference, presents significantly greater challenges than its matrix counterpart.
The contributions and novelty of our work are summarized as follows.
@@
  • Theoretical Contributions: We establish statistical guarantees for our distributed methods, demonstrating that they achieve a sharp statistical error rate that aligns with the minimax optimal rate possible in a non-distributed setting. Additionally, we calculate the asymptotic distribution of the proposed distributed estimator. This calculation aids in statistical inference and supports the construction of confidence regions for the singular subspaces. These contributions represent a theoretical advancement in utilizing aggregated common components, a finding that has not been previously documented in tensor analysis literature and thus enhances the existing computational approaches to distributed Tucker decomposition.
@@ 1.1 Related works
Tensor decompositions have become increasingly important in machine learning, electrical engineering, and statistics. While low-rank matrix decomposition theory is well-established, tensors present unique challenges, with multiple notions of low-rankness (Kolda and Bader, 2009), primarily CANDECOMP/PARAFAC (CP) and multilinear/Tucker decompositions. Early work by Richard and Montanari (2014), Hopkins et al. (2015), and Perry et al. (2020) focused on rank-1 spiked tensor PCA models, proposing methods like tensor unfolding and sum of square optimization. Zhang and Xia (2018) later developed a general tensor SVD framework based on Tucker decomposition, with Xia et al. (2022) extending this to inference for tensor PCA. Recent advances include sparse tensor SVD (Zhang and Han, 2019), binary tensor decomposition (Wang and Li, 2020), semi-parametric tensor factor models with covariates (Chen et al., 2024), and theoretical guarantees for Tucker/CP decomposition with correlated entries and missing data (Chen et al., 2024a, b).
Part of our analysis under the homogeneous scenario utilizes the decomposition technique in Xia et al. (2022) established for a non-distributed environment. Compared to their work, we highlight several major innovations in our study from the theoretical perspective. In the homogeneous setting, we conduct a delicate bias-variance analysis for the error decomposition to establish the theoretical results for the distributed estimator. For the heterogeneous setting, we develop the decomposition for the common-component estimators using a different technique from Xia et al. (2022), due to the existence of distinct individual components. Additionally, the procedure for determining the common component ranks and the extension to data-adaptive weighted estimators in the knowledge transfer setting are, to our best knowledge, new to the literature.
@@
To handle the challenges posed by massive and decentralized data, there has been a significant amount of recent literature developing distributed estimation or inference techniques for a variety of statistical problems (Li et al., 2013; Chen and Xie, 2014; Garber et al., 2017; Lee et al., 2017; Shi et al., 2018; Jordan et al., 2019; Fan et al., 2019; Volgushev et al., 2019; Chen et al., 2019; Li et al., 2020; Yu et al., 2020; Tan et al., 2022; Luo et al., 2022; Yu et al., 2022; Chen et al., 2022, 2024). We refer readers to Gao et al. (2022) for a comprehensive review. Among these works, our paper is most closely related to the literature on distributed matrix PCA. Fan et al. (2019) notably proposed a convenient one-shot approach that computes the top-$K$-dim eigenspace of the covariance matrix on each local machine and aggregates them on a central machine. They established a rigorous statistical guarantee showing that this approach achieves a sharp error rate as long as the sample size on each local machine is sufficiently large. An alternative multi-round method was simultaneously developed by Garber et al. (2017), which mainly estimates the first eigenspace by leveraging shift-and-invert preconditioning. Subsequently, Chen et al. (2022) proposed an improved multi-round algorithm for estimating the top-$K$-dim eigenspace that enjoys a fast convergence rate under weaker restrictions compared to Fan et al. (2019). Other works focused on providing more general statistical error bounds (Zheng and Tang, 2022) or improving the communication efficiency (Huang et al., 2021; Charisopoulos et al., 2021).
@@
The structure of this paper is as follows: Section 2 introduces our distributed tensor PCA algorithm for homogeneous settings and establishes its statistical error rates. Section 3 adapts the algorithm for heterogeneous environments and includes theoretical analysis. Section 4 explores knowledge transferring from source tensors to a target tensor within a heterogeneous setting. Section 5 presents simulations and real data analyses to evaluate the performance of our methods. Some commonly used notations in this paper are defined in Section A of the supplementary material. Sections B and C provide supplemental theoretical results regarding Theorems 2.1 and 4.1, respectively. Section E, notably, presents the asymptotic distribution of our proposed estimator, featuring distributed inference. All technical proofs and additional numerical experiments are presented in Sections D and F of the supplementary material, respectively.
@@ 2 Estimation for Homogeneous Tensors
In this section, we present the distributed tensor PCA algorithm for the homogeneous setting where each machine observes a tensor generated from the same underlying model. We first formulate the problem setup and illustrate the pooled estimator as a benchmark in Section 2.1, followed by our estimator and the intuition behind its construction detailed in Section 2.2. Finally, Section 2.3 establishes theoretical guarantees on the statistical error rate for our estimator.
@@ 2.1 Problem Setup and the Pooled Estimator
If the tensors $\{\mathcal{T}_{\ell}\}$ were allowed to be pooled onto a central machine, a standard way for estimating the singular space $\mathrm{Col}(U_{j})$ would be to conduct a Tucker decomposition on the averaged tensor $\overline{\mathcal{T}}=\frac{1}{L}\sum\limits_{\ell=1}^{L}\mathcal{T}_{\ell}=\mathcal{T}^{*}+\overline{\mathcal{Z}}$, where $\overline{\mathcal{Z}}=\frac{1}{L}\sum\limits_{\ell=1}^{L}\mathcal{Z}_{\ell}$ is a tensor with i.i.d. $\mathcal{N}(0,\frac{\sigma^{2}}{L})$ entries. Denote the estimator obtained through this method as $\widehat{U}_{\mathrm{pooled},j}$, $j\in[J]$, which serves as a benchmark against which we evaluate the effectiveness of our approach in the distributed environment.
@@ 2.3 Theoretical Guarantee
Assume that there exist constants $C_{1}$, $c_{1}$, and $C_{2}$ such that, with probability at least $1-C_{1}e^{-c_{1}p}$, $\big{\|}\widehat{U}_{j,\ell}^{(0)}\widehat{U}_{j,\ell}^{(0)\top}-U_{j}U_{j}^{\top}\big{\|}_{2}\leq C_{2}\sqrt{p}\sigma\lambda_{\min}^{-1}$ for any $j\in[J]$ and $\ell\in[L]$. If $p_{j}\asymp p$ for all $j$, $L=O(p^{c_{3}})$ for some $c_{3}>0$, $r^{J-1}=O(p)$, $\kappa_{0}=O(1)$, and $\sqrt{pr}=o(\lambda_{\min}/\sigma)$, then we have
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}\leq\widetilde{C}_{2}\bigg{(}\frac{\sigma}{\lambda_{\min}}\sqrt{\frac{pr}{L}}+\frac{pr\sigma^{2}}{\lambda_{\min}^{2}}\bigg{)},$ (3)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$ for some constants $\widetilde{C}_{1}$ , $\widetilde{c}_{1}$, and $\widetilde{C}_{2}$, where $\{\widehat{U}_{j}\}$ is the output of Algorithm 1.
@@
where $G_{1}=\mathcal{M}_{1}(\mathcal{G})$, $Z_{1,\ell}=\mathcal{M}_{1}(\mathcal{Z}_{\ell})$, $U_{1\perp}$ is the orthogonal complement of $U_{1}$, and $R_{1,\ell}$ is a remainder term. When the signal-to-noise ratio (SNR) satisfies $\sqrt{pr}=o(\lambda_{\min}/\sigma)$, the Frobenius norm of the first two mean-zero terms on the RHS of (4) is of the order $O\left(\sqrt{pr}\sigma\lambda_{\min}^{-1}\right)$, and the remainder term $R_{1,\ell}$ has a higher order $O\big{(}pr\sigma^{2}\lambda_{\min}^{-2}\big{)}$. By averaging the projection matrices on all machines, the order of the first two terms can be reduced to $O\big{(}\sqrt{pr}\sigma\lambda_{\min}^{-1}L^{-1/2}\big{)}$, while that of $R_{1,\ell}$ does not change, leading to the error rate in (3). When the SNR is sufficiently large with respect to $L$, the first term in (3) dominates the second one, which leads to the following corollary.
@@
$\sup_{j}\rho\big{(}\widehat{U}_{j},U_{j}\big{)}\leq\frac{\widetilde{C}_{2}\sigma}{\lambda_{\min}}\sqrt{\frac{pr}{L}},$ (5)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$ for some constants $\widetilde{C}_{1}$ , $\widetilde{c}_{1}$, and $\widetilde{C}_{2}$.
@@
Remark 3.
The second-order term $pr\sigma^{2}\lambda_{\min}^{-2}$ in Theorem 2.1 cannot be improved. We establish this through a lower bound result in Section B of the supplementary material, which shows that when the condition $\lambda_{\min}/\sigma\gtrsim\sqrt{prL}$ is violated, there exists a target tensor $\mathcal{T}^{*}$ and a set of initial estimators that, despite satisfy all the assumptions in Theorem 2.1, achieve only the sub-optimal convergence rate $pr\sigma^{2}\lambda_{\min}^{-2}$. This result demonstrates the sharpness of the error rate in (3) and implies the necessity of the condition $\lambda_{\min}/\sigma\gtrsim\sqrt{prL}$ in Corollary 2.1 for achieving the minimax optimal rate.
@@ 3 Estimation for Heterogeneous Tensors
$\mathcal{T}_{\ell}^{*}=\mathcal{G}_{\ell}\times_{1}[U_{1}\;V_{1,\ell}]\times_{2}[U_{2}\;V_{2,\ell}]\cdots\times_{J}[U_{J}\;V_{J,\ell}],$ (7)
where $U_{j}\in\mathbb{O}^{p_{j}\times r_{j,U}}$, $V_{j,\ell}\in\mathbb{O}^{p_{j}\times r_{j,V,\ell}}$, $U_{j}^{\top}V_{j,\ell}=0$, $\mathcal{G}_{\ell}\in\mathbb{R}^{r_{1,\ell}\times r_{2,\ell}\cdots\times r_{J,\ell}}$, and $r_{j,\ell}=r_{j,U}+r_{j,V,\ell}$. For each $j$, the component $U_{j}$ spans a common singular subspace shared by all tensors $\{\mathcal{T}_{\ell}\}$, while $V_{j,\ell}$ spans an individual subspace specific to each tensor. Moreover, the core $\mathcal{G}_{\ell}$ is allowed to be different for different $\ell$. The goal is to estimate the shared singular subspace and the individual subspace separately for $j\in[J]$ and $\ell\in[L]$. Note that in this section, we assume the noise level $\sigma^{2}$ is identical on all machines for clear presentation. The heterogeneity on $\sigma^{2}$ will be further investigated in Section 4.
@@
where $\Lambda_{j,\ell}$ is a diagonal matrix with decreasing singular values for all $j\in[J]$, $\ell\in[L]$.
@@
Condition (8) is closely related to the all-orthogonality condition in the HOSVD framework (De Lathauwer et al., 2000a). The all-orthogonality condition states that the mode-$j$ subtensors of the core tensor (i.e., the subtensor obtained by fixing the $j$-th index and letting the other indices be free) are mutually orthogonal with respect to the scalar product, which leads to orthogonal rows in $\mathcal{M}_{j}(\mathcal{G}_{\ell})$ and implies (8).
To clearly illustrate how this condition guarantees the uniqueness of $\mathrm{Col}(U_{j})$, we first restate the non-identifiablility of our heterogeneous model, $\mathcal{T}^{*}_{\ell}=\mathcal{G}_{\ell}\times_{1}[U_{1,\ell}\,V_{1,\ell}]\times_{2}[U_{2,\ell}\,V_{2,\ell}]\cdots\times_{J}[U_{J,\ell}\,V_{J,\ell}]$, where we initially do not assume that the components $U_{j,\ell}$ are shared across the tensors. This model is non-identifiable: it remains unchanged under the transformation $\widetilde{\mathcal{G}}_{\ell}=\mathcal{G}_{\ell}\times_{1}O^{\top}_{1,\ell}\times_{2}O^{\top}_{2,\ell}\cdots\times_{J}O^{\top}_{J,\ell}$ and $\big{[}\widetilde{U}_{j,\ell}\,\widetilde{V}_{j,\ell}\big{]}=[U_{j,\ell}V_{j,\ell}]O_{j,\ell}$, for any orthogonal matrices $\{O_{j,\ell}\in\mathbb{O}^{r_{j,\ell}\times r_{j,\ell}}\}_{j\in[J]}$. Note that $\mathcal{M}_{j}\big{(}\mathcal{G}_{\ell}\big{)}=O_{j,\ell}\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}_{\ell}\big{)}\big{(}\bigotimes\limits_{j^{\prime}\neq j}O_{j^{\prime},\ell}^{\top}\big{)}$. There exists a unique $O_{j,\ell}$—the left singular vector matrix of $\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}_{\ell}\big{)}$—such that $\mathcal{M}_{j}(\mathcal{G}_{\ell})\mathcal{M}_{j}(\mathcal{G}_{\ell})^{\top}=O_{j,\ell}^{\top}\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}_{\ell}\big{)}\mathcal{M}_{j}\big{(}\widetilde{\mathcal{G}}_{\ell}\big{)}^{\top}O_{j,\ell}$ is a diagonal matrix with decreasing singular values. This uniquely determines $[U_{j,\ell}\,V_{j,\ell}]=\big{[}\widetilde{U}_{j,\ell}\,\widetilde{V}_{j,\ell}\big{]}O_{j,\ell}^{\top}$. Based on this condition, we further assume that $U_{j,\ell}=U_{j}$ to be the common components shared across all tensors, yielding model (7).